Edna has found your dream vehicle that it is priced at $30,000. She can arrange for...

  1. Edna has found your dream vehicle that it is priced at $30,000. She can arrange for financingfrom the bank with 0 down and monthly payments (at the end of each month) at an annualized interest rate of 12% (r = 0.12), with a term of 60 months. If she buys it, it is hers. Alternatively, the dealer offers to lease it to her for four years. The terms of the lease are monthly payments (end of month) of $400 for 48 months plus a down payment of $4,000. At the end of the lease the car ownership reverts to the dealer.
    1. If she finances the car, what is Edna’s monthly payment? Hint: first calculate Edna’smonthly interest rate, rm; then find her monthly payment. Remember that her first payment doesn’t occur until the end of the first month!
    2. Find the market value of the car at the end of year four which would make Edna indifferent between leasing and buying.
Answer #1

a.

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 30,000, N = 60, FV = 0 ,I = 0.12/12]

PMT = $667.33

Monthly Payment = $667.33

b.

Cost of Car = $30,000

Downpayment = $4,000

Effective leased = 30,000 - 4,000 = $26,000

Calculating Value after 4 years,

Using TVM Calculation,

FV = [PV = 26,000, PMT = -400, N = 48, I = 0.12/12]

FV = $17,428.84

Value of car after 4 years = $17,428.84